# Force-current structure in Markovian open quantum systems and its applications: geometric housekeeping-excess decomposition and thermodynamic trade-off relations

### v1

@@ I Introduction
The thermodynamic theory of nonequilibrium systems has been explored since the last century, both in classical and quantum systems [1, 2]. The significant development of stochastic thermodynamics in this century [3, 4] has been pushing forward our understanding of not only classical but also quantum systems [5, 6]. The second law of thermodynamics is believed not to be violated even in quantum mechanics, and extensions of modern thermodynamic laws unveiled by stochastic thermodynamics, such as the fluctuation theorem [7] or the thermodynamic uncertainty relations [8], to quantum systems have been actively studied [5, 9, 10].
However, there is a structure that has been underexplored in quantum thermodynamics, i.e., the structure of irreversible currents and thermodynamic forces [11, 1, 12]. In classical systems, from mesoscopic stochastic systems to macroscopic nonlinear systems, it is known that the entropy production rate $\dot{\Sigma}$ is provided as the product between currents $J_{\alpha}$ and forces $F_{\alpha}$ as
@@
Although this is regarded as a fundamental law of classical nonequilibrium thermodynamics, there are few studies on quantum extension. An earlier form of the quantum extension of the force-current structure can be found in papers such as Refs. [13, 14], which are based on the continuous measurement. We can regard Ref. [15] as a sophistication of their idea, which relies on the eigenbasis of the density operator. While their approach captures an aspect of the profound connection between quantum and classical nonequilibrium thermodynamics, it still has some problems. First, since the expression relies on a specific basis, it lacks not only formal elegance but also theoretical transparency. We may be unable to identify what is quantum and what is not. In addition, as discussed in the main text, the formulation can encounter a problem when defining conservative and nonconservative forces. This is crucial because, in short, this separation of force corresponds to the separation between equilibrium and nonequilibrium systems.
@@
Further, we apply the developed framework to extend results obtained for classical systems in recent years. We first derive the geometric housekeeping-excess decomposition of the entropy production rate [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. The housekeeping-excess decomposition [18] splits the total dissipation into one part to keep the system out of equilibrium (the housekeeping dissipation) and the other stemming from the system’s dynamics (the excess dissipation). As it can tighten thermodynamic inequalities such as the second law, it enables us to obtain a more precise understanding of irreversibility and fundamental limitations in the dynamics [18, 19, 25, 36, 37, 26, 27, 28, 29, 32, 30, 31]. The geometric approach [32], which we adopt, does not require any particular steady state, unlike the conventional one [34, 35], so it provides a more general way of decomposition.
In addition, we show that the framework reveals thermodynamic trade-off relations in quantum dynamics. Thermodynamic trade-off relations such as the thermodynamic uncertainty relations [8, 38, 39, 40, 41, 42, 9, 43] or the thermodynamic speed limits [36, 37, 44, 45, 26, 27, 28, 29, 43], involving the entropy production, provide the dynamics with fundamental limitations tighter than the second law of thermodynamics. We derive trade-off relations between the entropy production rate and fluctuations or a typical time scale from the force-current structure. The former one is the quantum extension of the short-time thermodynamic uncertainty relation [46]. We will explain that this generic inequality leads to some specific results reported recently [15, 47]. The latter is similar to the dissipation-time uncertainty relation [48] and the time-energy uncertainty relation [49]. In discussing these trade-offs, we will introduce a measure of fluctuation, which we term the quantum diffusivity. We show that the quantum diffusivity coincides with a classical counterpart when the observable is classical.
@@ II Preliminary
II.1 Open quantum system
@@
Here, $H\in\mathrm{herm}(\mathcal{H})$ is the Hamiltonian, the reduced Planck constant $\hbar$ is set to one, and $K_{\mathrm{all}}$ is the set of jumps. Each $k\in K_{\mathrm{all}}$ has a positive rate $\gamma_{k}>0$ and a jump operator $L_{k}\in\mathrm{opr}(\mathcal{H})$, which represents state changes. We assume that every $k\in K_{\mathrm{all}}$ has a unique counterpart $-k\in K_{\mathrm{all}}$ such that $L_{-k}=L_{k}^{\dagger}$. Let $K$ be the set of indexes that only includes either of each pair. For later convenience, we define for each $k\in K$
@@
Thermodynamics is introduced to the model by imposing the local detailed balance condition $\ln(\gamma_{k}/\gamma_{-k})=s_{k}$ for each pair $(k,-k)$, where $s_{k}$ is the entropy flux into the environment ($s_{-k}=-s_{k}$[14]. The entropy production rate (EPR) of the total system (system + environment) is given as
@@
The first term gives the von Neumann entropy change $-(d/dt)k_{\mathrm{B}}\mathop{\mathrm{tr}}\nolimits(\rho\ln\rho)$, and the second provides the entropy change in the environment. Hereafter, we set the Boltzmann constant $k_{\mathrm{B}}$ to one for simplicity.
@@
where $\omega_{k}$ is an energy gap corresponding to $k$ and $\beta_{k}\omega_{k}$ provides the entropy change in the bath. In addition, if $\beta_{k}=\beta$ for every $k$ with a single parameter $\beta$ holds, the system can have an equilibrium state $\rho^{\mathrm{eq}}=e^{-\beta H}/Z$ with $Z=\mathop{\mathrm{tr}}\nolimits(e^{-\beta H})$, where the EPR vanishes. We call systems satisfying the conditions in Eq. (6) with $\beta_{k}=\beta$ detailed balanced.
Throughout the paper, we always assume the Hamiltonian, the jump operates, and the rates are time-independent; namely, we only consider autonomous systems. However, this does not lose the generality of the results, and they can easily be extended to non-autonomous systems.
@@ II.2 Classical expression of the quantum master equation
These classical expressions have been utilized in exploring quantum thermodynamics (e.g., [15, 44, 43]); however, we can find some unnatural points in this method. First, the transition rates change in time, even if $\gamma_{k}$ and $L_{k}$ do not because the eigenbasis $|i\rangle$ is time-dependent. Second, the current and force have nothing to do with any operator expression, so the connection to quantum mechanics is lost. The first point can be attributed to the Hamiltonian dynamics, which should not be directly associated with the dissipation because it is supposed to govern the reversible dynamics. In fact, our definition will never refer to the Hamiltonian. The second problem is also resolved in our formulation; we will introduce operators expressing currents and forces. We will provide a method to treat quantum thermodynamics as a quantum theory without relying on any basis. In addition to these problems, another one exists in the conventional definition of force and current, as discussed in the next section.
@@ II.3 Brief review of classical systems
We can equip the dynamics with a thermodynamic structure by assuming that the currents are split into forward and backward fluxes $J_{e}^{+}>0$ and $J_{e}^{-}>0$, as $J_{e}=J_{e}^{+}-J_{e}^{-}$ and the local detailed balance $F_{e}=\ln(J_{e}^{+}/J_{e}^{-})$. Here, $F_{e}$ is the thermodynamic force on process $e$ and provides the EPR $\dot{\Sigma}=\bm{J}^{\mathrm{T}}\bm{F}=\sum_{e\in E}J_{e}F_{e}$. Let us examine these quantities in a little more depth.
@@
Second, the EPR has an additional geometric interpretation, whose utility will be shown in later sections. We define a diagonal matrix $\mathsf{M}=(M_{e}\delta_{ee^{\prime}})$ by $M_{e}=(J_{e}^{+}-J_{e}^{-})/\ln(J_{e}^{+}/J_{e}^{-})$ ($\delta_{ee^{\prime}}$ is the Kronecker delta), which connects the force and current as $\bm{J}=\mathsf{M}\bm{F}$. The diagonal elements of $\mathsf{M}$ are the so-called logarithmic means between $\bm{J}^{+}$ and $\bm{J}^{-}$ [52]. Since the log mean is always positive as long as the two numbers are positive, $\mathsf{M}$ is a positive definite matrix. Therefore, the EPR can be seen as the squared norm of $\bm{F}$ with $\mathsf{M}$ being the metric, $\dot{\Sigma}=\bm{F}^{\mathrm{T}}\mathsf{M}\bm{F}$. Consequently, we can decompose the EPR into the housekeeping and excess parts in general by projecting the force onto the conservative subspace orthogonally regarding this metric (if interested, see also [26, 27, 33, 29, 31]).
@@
III Reformulate quantum thermodynamics
@@ III.1 Current operator and force operator
where $\otimes$ means the tensor product and $\oplus$ the direct sum of operators. These operators are supposed to separately represent the transition structure and the transition rates in the dynamics. Hereafter, we do not distinguish the number zero and the zero operators and write them as 00.
Then, regarding $\Gamma\otimes\rho$ as the fluxes ($\bm{J}^{\pm}$) in the classical case, we introduce the current and the force operator as follows:
@@
We comment here that an expression similar to Eq. (13) is provided recently in Ref. [53].
@@ III.2 Connection to dynamics
When the system is detailed balanced, the system has an equilibrium solution $\rho^{\mathrm{eq}}=e^{-\beta H}/Z$, where the current vanishes $\mathbb{J}(\rho^{\mathrm{eq}})=0$. Note that, in a general steady state $\rho^{\mathrm{ss}}$, we only have $-i[H,\rho^{\mathrm{ss}}]+\nabla_{\mathbb{L}}^{*}\mathbb{J}(\rho^{\mathrm{ss}})=0$ and the current $\mathbb{J}(\rho^{\mathrm{ss}})$ can be nonzero. To show the statement, first note that the commutation relation $[L_{k},H]=\omega_{k}L_{k}$ lead to $L_{k}H^{n}=(H+\omega_{k})^{n}L_{k}$, which further brings $L_{k}e^{-\beta H}=e^{-\beta(H+\omega_{k})}L_{k}$  [^1] The first equality is proved by the mathematical induction; $L_{k}H=(H+\omega_{k})L_{k}$ holds and assuming $L_{k}H^{n}=(H+\omega_{k})^{n}L_{k}$ leads to $L_{k}H^{n+1}=(H+\omega_{k})^{n}L_{k}H=(H+\omega_{k})^{n+1}L_{k}$. The second one is derived by applying this equality to each term of the Taylor expansion of $e^{-\beta H}$. . Applying this to Eq. (13), we get

### v3

@@ I Introduction
The thermodynamic theory of nonequilibrium systems has been explored since the last century, both in classical and quantum systems [1, 2, 3]. The significant development of stochastic thermodynamics in this century [4, 5] has been pushing forward our understanding of not only classical but also quantum systems [6, 7, 3]. The second law of thermodynamics is believed not to be violated even in quantum mechanics, and extensions of modern thermodynamic laws unveiled by stochastic thermodynamics, such as the fluctuation theorem [8] or the thermodynamic uncertainty relations [9], to quantum systems have been actively studied [6, 10, 11].
However, there is a structure that has been underexplored in quantum thermodynamics, i.e., the structure of irreversible currents and thermodynamic forces [12, 13, 1, 14]. In classical systems, from mesoscopic stochastic systems to macroscopic nonlinear systems, it is known that the entropy production rate $\dot{\Sigma}$ is provided as the product between currents $J_{\alpha}$ and forces $F_{\alpha}$ as
@@
Although this is regarded as a fundamental law of classical nonequilibrium thermodynamics, there are few studies on quantum extensions. One important methodology was established in Ref. [15] and similar results have repeatedly been presented in the literature [16, 17, 18]. While their approach captures an aspect of the profound connection between quantum and classical nonequilibrium thermodynamics, it still has some problems. First, since the expression relies on a specific basis, it lacks not only formal elegance but also theoretical transparency. We may be unable to identify what is quantum and what is not. In addition, as discussed in the main text, the formulation can encounter a problem when defining conservative and nonconservative forces. This is crucial because, in short, this decomposition of forces corresponds to the separation between equilibrium and nonequilibrium systems.
@@
Further, we apply the developed framework to extend results obtained for classical systems in recent years. We first derive the geometric housekeeping-excess decomposition of the entropy production rate [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. The housekeeping-excess decomposition [21] splits the total dissipation into one part to keep the system out of equilibrium (the housekeeping dissipation) and the other stemming from the system’s dynamics (the excess dissipation). As it can tighten thermodynamic inequalities such as the second law, it enables us to obtain a more precise understanding of irreversibility and fundamental limitations in the dynamics [21, 22, 28, 40, 41, 29, 30, 31, 32, 33, 36, 34, 35]. The geometric approach [36], which we adopt, does not require any particular steady state, unlike the conventional one [38, 39], so it provides a more general way of decomposition.
In addition, we show that the framework reveals thermodynamic trade-off relations in quantum dynamics. Thermodynamic trade-off relations such as the thermodynamic uncertainty relations [9, 42, 43, 44, 45, 46, 10, 47] or the thermodynamic speed limits [40, 41, 48, 49, 29, 30, 31, 32, 33, 47], involving the entropy production, provide the dynamics with fundamental limitations tighter than the second law of thermodynamics. We derive trade-off relations between the entropy production rate and fluctuations or a typical time scale from the force-current structure. The former one is the quantum extension of the short-time thermodynamic uncertainty relation [50]. We will explain that this generic inequality leads to some specific results reported recently [18, 51]. The latter is similar to the dissipation-time uncertainty relation [52] and the time-energy uncertainty relation [53]. In discussing these trade-offs, we will introduce a measure of fluctuation, which we term the quantum diffusivity. We show that the quantum diffusivity coincides with a classical counterpart when the observable is classical.
@@ II Preliminary
II.1 Open quantum systems
@@
[…]
@@
Thermodynamics is introduced to the model by imposing the local detailed balance condition $\ln(\gamma_{k}/\gamma_{-k})=s_{k}$ for each pair $(k,-k)$, where $s_{k}$ is the entropy flux into the environment ($s_{-k}=-s_{k}$[17]. The entropy production rate (EPR) of the total system (system + environment) is given as [3]
@@
The first term gives the von Neumann entropy change $-(d/dt)k_{\mathrm{B}}\mathop{\mathrm{tr}}\nolimits(\rho\ln\rho)$, and the second provides the entropy change in the environment. As shown later, the EPR is always non-negative owing to the local detailed balance condition. Hereafter, we set the Boltzmann constant $k_{\mathrm{B}}$ to one for simplicity.
@@
where $\omega_{k}$ is the energy gap corresponding to $k$ and $\beta_{k}\omega_{k}$ provides the entropy change $s_{k}$ in the bath. In this case, the second term in the right-hand side of Eq. (5) is given by $-\sum_{k}\beta_{k}\mathop{\mathrm{tr}}\nolimits(H\mathcal{D}_{k}(\rho))$ and $-\mathop{\mathrm{tr}}\nolimits(H\mathcal{D}_{k}(\rho))$ is interpreted as the heat flux into bath $k$ [3]. In addition, if $\beta_{k}=\beta$ for every $k$ with a single parameter $\beta$ holds, the system can have an equilibrium state $\rho^{\mathrm{eq}}=e^{-\beta H}/Z$ with $Z=\mathop{\mathrm{tr}}\nolimits(e^{-\beta H})$, where the EPR vanishes. We call systems satisfying the conditions in Eq. (6) with $\beta_{k}=\beta$ detailed balanced.
Throughout the paper, we always assume the Hamiltonian, the jump operators, and the rates are time-independent; namely, we only consider autonomous systems. However, this does not restrict the generality of the results, and they can easily be extended to non-autonomous systems because we only deal with quantities defined at each time, and time derivatives are not generally involved.
@@ II.2 Classical expression of the quantum master equation
These classical expressions have been utilized in exploring quantum thermodynamics (e.g., [18, 48, 47]); however, we can find some issues in this method. First, the transition rates change in time, even if $\gamma_{k}$ and $L_{k}$ do not because the eigenbasis $|i\rangle$ is time-dependent. Second, the current and force have nothing to do with any operator expression, so the connection to quantum mechanics is lost. The first point can be attributed to the Hamiltonian dynamics, which should not be directly associated with the dissipation because it is supposed to govern the reversible dynamics. In fact, our definition will never refer to the Hamiltonian. The second problem is also resolved in our formulation; we will introduce operators expressing currents and forces. We will provide a method to treat quantum thermodynamics as a quantum theory without relying on any basis. In addition to these problems, another one exists in the conventional definition of force and current, as discussed in the next section.
@@ II.3 Brief review of classical systems
We can equip the dynamics with a thermodynamic structure by assuming that the currents are split into forward and backward fluxes $J_{e}^{+}>0$ and $J_{e}^{-}>0$, as $J_{e}=J_{e}^{+}-J_{e}^{-}$ and the local detailed balance $F_{e}=\ln(J_{e}^{+}/J_{e}^{-})$. Here, $F_{e}$ is the thermodynamic force on process $e$ and provides the EPR $\dot{\Sigma}=\bm{J}^{\mathrm{T}}\bm{F}=\sum_{e\in E}J_{e}F_{e}$, where the EPR is in advance defined by the system’s entropy change and environmental entropy increase [1, 4]. Let us examine these quantities in a little more depth.
@@
Second, the EPR has an additional geometric interpretation, whose utility will be shown in later sections. We define a diagonal matrix $\mathsf{M}=(M_{e}\delta_{ee^{\prime}})$ by $M_{e}=(J_{e}^{+}-J_{e}^{-})/\ln(J_{e}^{+}/J_{e}^{-})$ ($\delta_{ee^{\prime}}$ is the Kronecker delta), which connects the force and current as $\bm{J}=\mathsf{M}\bm{F}$. The diagonal elements of $\mathsf{M}$ are the so-called logarithmic means between $\bm{J}^{+}$ and $\bm{J}^{-}$ [56]. Since the log mean is always positive as long as the two numbers are positive, $\mathsf{M}$ is a positive definite matrix. Therefore, the EPR can be seen as the squared norm of $\bm{F}$ with $\mathsf{M}$ being the metric, $\dot{\Sigma}=\bm{F}^{\mathrm{T}}\mathsf{M}\bm{F}$. Consequently, we can decompose the EPR into the housekeeping and excess parts in general by projecting the force onto the conservative subspace orthogonally regarding this metric (if interested, see also [29, 30, 37, 32, 33, 35]).
@@
III Reformulating quantum thermodynamics
@@ III.1 Current operator and force operator
where $\otimes$ means the tensor product and $\oplus$ the direct sum of operators (the direct sum can be seen as a diagonal alignment of the blocks of operators/real numbers). These operators are supposed to separately represent the transition structure and the transition rates in the dynamics. Hereafter, we do not distinguish the number zero and the zero operators and write them as 00.
The introduction of an auxiliary space is essential; in classical systems, a jump can be labeled by a pair of states $i,j$ without introducing a special label for jumps. On the other hand, in quantum dynamics, a jump can involve more than two (eigen)states. In our formulation, an auxiliary space is necessary to treat quantum jumps apart from specific states, and it enables us to discuss connections between dynamics and thermodynamics more efficiently. As will be discussed later, the characterization of conservative systems in terms of thermodynamic forces and the definition of the geometric housekeeping-excess decomposition are only possible with the introduction of an auxiliary space.
Next, regarding $\Gamma\otimes\rho$ as the fluxes ($\bm{J}^{\pm}$) in the classical case, we introduce the current and the force operator as follows:
@@
The reason we regard these operators as currents and forces is that they satisfy the dynamical and thermodynamic relationships that classical currents and forces fulfill, as will be shown in the following. A few of them are listed in Table 1. Some correspondence to the classical notion is also suggested in Example V.1. We comment here that an expression similar to Eq. (13) is provided recently in Ref. [57].
@@
Now that the three properties of $\mathcal{M}_{\Gamma\otimes\rho}$, linearity, symmetry, and positive definiteness, are proved, we can connect Eq. (16) to the renowned Onsager reciprocal relation [12, 13]. Let us consider physical currents, e.g., electric currents or heat currents, $j_{\alpha}(\rho)\in\mathbb{R}$ ($\alpha=1,2,\dots$) given by
$j_{\alpha}(\rho)=\langle\mathbb{A}_{\alpha},\mathbb{J}(\rho)\rangle$ (22)
with anti-Hermitian operators $\mathbb{A}_{\alpha}\in\mathrm{anti}(\mathfrak{h}\otimes\mathcal{H})$. We assume that the force operator is given in the dual form
$\mathbb{F}(\rho)=\sum_{\alpha}f_{\alpha}(\rho)\mathbb{A}_{\alpha}$ (23)
by physical forces $f_{\alpha}(\rho)\in\mathbb{R}$, such as voltage or temperature gradient. We further define a matrix $L(\rho)$ by
$L_{\alpha\beta}(\rho)\coloneqq\langle\mathbb{A}_{\alpha},\mathbb{A}_{\beta}\rangle_{\Gamma\otimes\rho}=\langle\mathbb{A}_{\alpha},\mathcal{M}_{\Gamma\otimes\rho}(\mathbb{A}_{\beta})\rangle.$ (24)
If $\rho=\rho^{\mathrm{eq}}=e^{-\beta H}/Z$, this definition resembles the Green–Kubo formula [59, 60] as
$L_{\alpha\beta}(\rho^{\mathrm{eq}})$
$=\frac{1}{Z}\int_{0}^{1}\mathop{\mathrm{tr}}\nolimits[\mathbb{A}_{\alpha}(\Gamma^{s}\otimes e^{-s\beta H})\mathbb{A}_{\beta}(\Gamma^{1-s}\otimes e^{-(1-s)\beta H})]ds.$ (25)
Then, Eq. (16) leads to the relation
$j_{\alpha}(\rho)=\sum_{\beta}L_{\alpha\beta}(\rho)f_{\beta}(\rho)$ (26)
because $\sum_{\beta}L_{\alpha\beta}(\rho)f_{\beta}(\rho)=\sum_{\beta}\langle\mathbb{A}_{\alpha},\mathcal{M}_{\Gamma\otimes\rho}(f_{\beta}(\rho)\mathbb{A}_{\beta})\rangle=\langle\mathbb{A}_{\alpha},\mathcal{M}_{\Gamma\otimes\rho}(\mathbb{F}(\rho))\rangle$ and $\mathbb{J}(\rho)=\mathcal{M}_{\Gamma\otimes\rho}(\mathbb{F}(\rho))$. Let us discuss that Eq. (26) is a nonequilibrium extension of the Onsager reciprocal relation. When $\rho$ is close to equilibrium $\rho^{\mathrm{eq}}$, $j_{\alpha}(\rho)$ and $f_{\beta}(\rho)$ will be of the order of the difference between $\rho$ and $\rho^{\mathrm{eq}}$, which we write $\varepsilon$. Then, $L(\rho)$ can be replaced by $L(\rho^{\mathrm{eq}})$ because $[L_{\alpha\beta}(\rho)-L_{\alpha\beta}(\rho^{\mathrm{eq}})]f_{\beta}(\rho)=O(\varepsilon^{2})$. Thus, near the equilibrium state $\rho^{\mathrm{eq}}$, to leading order, we obtain
$j_{\alpha}(\rho)=\sum_{\beta}L_{\alpha\beta}(\rho^{\mathrm{eq}})f_{\beta}(\rho).$ (27)
Moreover, since $\mathbb{A}_{\alpha}$ are anti-Hermitian, the symmetry of the inner product $\langle\cdot,\cdot\rangle_{\Gamma\otimes\rho}$ proves the reciprocal relation
$L_{\alpha\beta}(\rho)=L_{\beta\alpha}(\rho).$ (28)
Therefore, Eq. (27) is the Onsager reciprocal relation [12, 13] and Eq. (26) is seen as its nonequilibrium extension [61]. If $\mathbb{A}_{\alpha}$ are linearly independent, $L(\rho)$ becomes positive definite because for any $\bm{\xi}=(\xi_{\alpha})$, we have
$\bm{\xi}^{\mathrm{T}}L(\rho)\bm{\xi}=\Big{\lVert}\sum_{\alpha}\xi_{\alpha}\mathbb{A}_{\alpha}\Big{\rVert}_{\Gamma\otimes\rho}^{2},$ (29)
and it is positive unless $\bm{\xi}=0$. In general, we have the following expression of the EPR
$\dot{\Sigma}(\rho)=\sum_{\alpha,\beta}f_{\alpha}(\rho)L_{\alpha\beta}(\rho)f_{\beta}(\rho).$ (30)
We finally mention that Eq. (23) is a crucial assumption and should be proved generally or verified in concrete cases.
@@ III.2 Connection to dynamics
When the system is detailed balanced, the system has an equilibrium solution $\rho^{\mathrm{eq}}=e^{-\beta H}/Z$, where the current vanishes $\mathbb{J}(\rho^{\mathrm{eq}})=0$. Note that, in a general steady state $\rho^{\mathrm{ss}}$, we only have $-i[H,\rho^{\mathrm{ss}}]+\nabla_{\mathbb{L}}^{*}\mathbb{J}(\rho^{\mathrm{ss}})=0$ and the current $\mathbb{J}(\rho^{\mathrm{ss}})$ can be nonzero. To show the statement, first note that the commutation relation $[L_{k},H]=\omega_{k}L_{k}$ leads to $L_{k}H^{n}=(H+\omega_{k})^{n}L_{k}$, which further brings $L_{k}e^{-\beta H}=e^{-\beta(H+\omega_{k})}L_{k}$  [^1] The first equality is proved by the mathematical induction; $L_{k}H=(H+\omega_{k})L_{k}$ holds and assuming $L_{k}H^{n}=(H+\omega_{k})^{n}L_{k}$ leads to $L_{k}H^{n+1}=(H+\omega_{k})^{n}L_{k}H=(H+\omega_{k})^{n+1}L_{k}$. The second one is derived by applying this equality to each term of the Taylor expansion of $e^{-\beta H}$. . Applying this to Eq. (13), we get
@@
In addition, the gradient operator gives a sufficient condition of conserved quantities in autonomous systems. If a time-independent observable $X$ satisfies $[H,X]=0$, the standard condition of conserved quantities in quantum mechanics, and
$\nabla_{\mathbb{L}}X=0,$ (39)
then we find $\langle X\rangle=\mathop{\mathrm{tr}}\nolimits(\rho X)$ is conserved because
$\frac{d}{dt}\langle X\rangle$ $=-i\mathop{\mathrm{tr}}\nolimits(X[H,\rho])+\mathop{\mathrm{tr}}\nolimits(X\mathcal{D}(\rho))$
$=i\mathop{\mathrm{tr}}\nolimits(\rho[H,X])+\langle X,\nabla_{\mathbb{L}}^{*}\mathbb{J}(\rho)\rangle$
$=i\mathop{\mathrm{tr}}\nolimits(\rho[H,X])+\langle\nabla_{\mathbb{L}}X,\mathbb{J}(\rho)\rangle=0.$
This is analogous to the situation in CRNs, where the stoichiometric matrix plays the same role as $\nabla_{\mathbb{L}}^{*}$ and no counterpart to the Hamiltonian exists [64, 27]. An example of such a conserved quantity can be found in the system discussed in Sec. V.2.